OpenStudy (anonymous):

AB&AC r tangent 2 d circle with centre osuch that

OpenStudy (anonymous):

140

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

ABOC forms a quadrilateral. 2 of the angles are 90. sum of the rest is 180. one is 40. so, the other is 140

OpenStudy (callisto):

SEE THE ATTACHMENT

OpenStudy (anonymous):

y u divide 40 with 2

OpenStudy (callisto):

because tangent properties tell you that you can divide the angle by 2when you join the angle point to the centre

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

plz answer my next question

OpenStudy (anonymous):

90+90=180+20=200-180=20????????????

OpenStudy (anonymous):

then hw its 140

OpenStudy (anonymous):

thanks bt i m still confuse????????